# Numerical simulation and analysis of mixing enhancement due to chaotic advection using an adaptive approach for approximating the dilution index

### v1

@@ Abstract
A velocity field characterized by chaotic advection induces stretching and folding processes that increase the solute-solvent interface available for diffusion. Identifying chaotic flow fields with optimized mixing enhancement is relevant for applications like groundwater remediation, microfluidics, and many more. This work uses the dilution index to quantify the temporal increase in mixing of a solute within its solvent. We introduce a new approach to select a suitable grid size for each time step in the dilution index approximation, motivated by the theory of representative elementary volumes. It preserves the central feature of the dilution index, which is monotonically increasing in time and hence leads to reliable results. Our analysis highlights the importance of a suitable choice for the grid size in the dilution index approximation. We use this approach to demonstrate the mixing enhancement for two chaotic injection-extraction systems that exhibit chaotic structures: a source-sink dipole and a rotated potential mixing. By analyzing the chaotic flow fields, we identify Kolmogorov–Arnold–Moser (KAM) islands, non-mixing regions that limit the chaotic area in the domain and, thereby, the mixing enhancement. Using our new approach, we assess the choice of design parameters of the injection-extraction systems to effectively engineer chaotic mixing. We demonstrate the important role of diffusion in filling the KAM islands and reaching complete mixing in the systems.
[…]
@@ 1 Introduction
The natural mixing processes in groundwater and microfluidic systems occurring by molecular diffusion are too slow for many applications Stone and Kim (2001); Stremler et al. (2004); Rolle and Le Borgne (2019). However, advection can boost mixing processes between solute and solvent. This is particularly evident when the flow leads to deformations of the plume, as with shear flow, focusing and defocusing flow, twisting flow Rolle and Le Borgne (2019); Paster et al. (2015), or helical flow Ye et al. (2015, 2016). It is important to note that while turbulent flows can enhance mixing Warhaft (2000), they are infeasible in certain situations, such as laminar flow in porous media. This limitation underscores the need for alternative methods. One of such alternatives is chaotic advection, a method that has shown potential to enhance fluid mixing Stremler et al. (2004); Liu et al. (2000); Aref (1984) and can also occur in flows at low Reynolds numbers Aref (1990). Due to the stretching and folding, the surface between solute and solvent increases, and concentration gradients steepen, boosting the diffusion appearing over this surface Ottino (1989); Borgne et al. (2014); Weeks and Sposito (1998). The process of engineering chaotic flow in a system, which requires increased mixing, is called "design for chaos" Stremler et al. (2004); Liu et al. (2000).
@@
Mixers within laminar flow regimes have a wide range of applications. Beginning with the topic of groundwater remediation, it is shown theoretically Lester et al. (2013), numerically Turuban et al. (2018) and experimentally Souzy et al. (2020); Heyman et al. (2020); Kree and Villermaux (2017) that the flow through porous media already exhibits chaotic trajectories at the pore scale. At the Darcy scale, the work of Bagtzoglou and Oates (2007) gives theoretical evidence that engineered chaotic advection using randomly oscillating wells enhances groundwater remediation if successfully implemented into groundwater flow systems. In practice, an implementation was achieved in Cho et al. (2019) where the rotated potential mixing (RPM) flow was engineered at field scale, leading to improved lateral tracer spreading. Another application is in microfluidics, where many processes are constrained by incomplete mixing in a laminar flow regime of low Reynolds numbers and high Péclet numbers Aref et al. (2017); Ward and Fan (2015). A range of passive mixing designs, such as flow-through mixers Stremler et al. (2004); Liu et al. (2000); Song et al. (2003) which are based on a given channel geometry, can be used to increase mixing involving sensitive species Aref et al. (2017). These mixers find application in the synthesis of biomolecules, drug delivery, and diagnostic testing Teh et al. (2008). On the contrary, by using active mixers, including microstirrers, acoustic mixers, and flow pulsation Ward and Fan (2015), chaotic advection is designed using external perturbation sources Aref et al. (2017). A successful implementation of active mixing using pulsating flow was achieved in Stremler et al. (2004) for enhancing gene expression profiling. Further applications of mixers range from food processing Metcalfe and Lester (2009) and heat transfer devices Aref et al. (2017); Ganesan et al. (1997) to the understanding of chaotic flow through blood vessels Schelin et al. (2009), showing that chaotic advection and the thereby increased degree of dilution is of high interest.
@@
Modeling transport processes requires solving the classical advection-diffusion equation Kitanidis (1994); Basilio Hazas et al. (2022). Besides the more classical Eulerian methods, like finite differences, finite elements, or finite volumes, one can approximate the solution to this equation using Lagrangian particle tracking. This approach is widely applied in the field of porous media for both conservative Tartakovsky et al. (2008); Le Borgne et al. (2008a, b) and reactive tracers Tartakovsky (2010); Paster et al. (2015). The Lagrangian perspective, a preferred method for chaotic advection setups, reveals more information about the fractal plume structure Bagtzoglou and Oates (2007) and is therefore used for this type of flow, among others, in Aref (2002); Bagtzoglou and Oates (2007). The passive advection of a particle is represented by the advection equation, an ordinary differential equation Aref (1990), while the diffusive process is added in form of a random walk Bagtzoglou and Oates (2007); Ghoniem and Sherman (1985). For two-dimensional incompressible flows, the advection equation is defined by a Hamiltonian system Bear (1988); Aref (1990, 1984); Chaté et al. (1999). The analysis of the resulting advective flow reveals basic properties. Of interest may be Lyapunov exponents, as they show the strength of the stretching and thereby quantify the extent of chaotic flow structures. For volume-preserving flow, we will also find fixed points that are either elliptic or hyperbolic. They alternate on a chain of periodic points and thereby form structures known as island chains Chaté et al. (1999). If the flow is chaotic, we may observe the occurrence of Kolmogorov–Arnold–Moser (KAM) islands, which are nonmixing regions arising around the elliptic points in the flow Lester et al. (2009); Chaté et al. (1999). In the remaining chaotic region, the flow is guided by homoclinic or heteroclinic tangles of the manifolds belonging to unstable stagnation points, leading to exponential fast deformations Chaté et al. (1999). Since KAM islands consist of periodic points, particles can only enter them by diffusion. Hence, they limit the chaotic area in the domain and decrease mixing Chaté et al. (1999).
@@
The goal of stretching and folding is to enlarge the interface between solute and solvent, thereby increasing diffusive fluxes over this interface Bagtzoglou and Oates (2007); Neupauer et al. (2014) and enhancing mixing processes. An approach to quantify the effect of mixing is the computation of the dilution index, an entropy-related approach to measure dilution Kitanidis (1994). The term dilution refers to the increase in volume of the fluid occupied by a solute. If the solute concentration follows the advection diffusion equation, this volume increases monotonically in time, hence one of the key properties of the dilution index is that it also increases monotonically in time Kitanidis (1994). For incompressible flow in a non-diffusive system, the volume of a conservative solute is constant in time, showing that the diffusion limits the increase in dilution. This measure, applied on results of particle tracking simulations, has a strong dependence on the grid size. For a non suitable grid the approximation of the dilution index may not show monotonic growth in time, as we see in Kapoor and Kitanidis (1998). We address this issue in one part of this paper.
@@ 2.1 Chaotic advection setups
In this paper, we consider flow in two dimensions. Hence, to successfully engineer chaos, the advection equation must be unsteady Ottino (1989); Stremler et al. (2004); Aref (1990). By following a large number of tracer particles inside the system, we approximate the spatial concentration distribution of the plume over time.
@@ 2.1.1 Pulsed source-sink system
The first setup under consideration is called the pulsed source-sink system introduced in Jones and Aref (1988) and later also addressed in Sposito and Weeks (1998); Stremler et al. (2004); Sposito (2006). We consider a two-dimensional plane with one sink and one source at arbitrary locations (see Fig. 1). During the experiment, the sink and the source are operated alternatively for a time interval of length $\tau$, meaning only one is switched on at a time. The particles extracted by the sink are reinjected by the source in the next stroke using the first-in-first-out procedure.
Figure 1: Two-dimensional plane with one sink and one source at ±$\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. All particles with a distance to the sink smaller than $\Lambda$ are extracted during the next stroke of the sink. Within the following stroke of the source, all of them are reinjected and occupy the circle around the source of radius $\Lambda$.
In general, the flow of an incompressible fluid in a non-deformable medium is described by a Lagrange stream function Bear (1988); Aref (1990, 1984); Chaté et al. (1999). The existence of a Lagrange stream function $\psi:\mathbb{R}^{s}\rightarrow\mathbb{R}$ implies that the associated flow can be described using the Hamiltonian system Sposito and Weeks (1998); Durst (2022)
@@
For a point source or sink at location $\left(\begin{smallmatrix}\bar{x}\\ \bar{y}\end{smallmatrix}\right)$, $\psi$ is given in Sposito (2006) by
@@
The analytical solution of the flow defined in (1) is derived in Sposito (2006); Jones and Aref (1988). We give a detailed derivation in Appendix A. Arguing that each positioning of sink and source can be transformed into an equivalent configuration with sink and source at ±$\pm\left(\begin{smallmatrix}a\\ 0\end{smallmatrix}\right)$ and by multiplying the length scale by $\frac{1}{a}$, we assume sink and source at locations ±$\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. The flow is parametrized by a combination of $\tau$, $Q$, and $a$ in the definition of
@@
It holds that all particles within the circle of radius $\Lambda$ around the sink are extracted during the operation of the sink for time $\tau$. This circle is depicted as $B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$ and also shown in Fig. 1. For the extraction-reinjection procedure, the authors of Jones and Aref (1988) derive the map $M_{\text{inj}}:B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\rightarrow\mathbb{R}^{2}$, taking into consideration the first-in-first-out reinjection protocol and an injection angle of $\pi-\theta$ given the extraction angle $\theta$. The map is given by
@@
$\begin{split}M\begin{pmatrix}x\\ y\end{pmatrix}=\begin{cases}\left(\Psi_{+}^{\tau}\circ\Psi_{-}^{\tau}\right)\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)&\text{for }\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in\mathbb{R}^{2}/B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\\ M_{\text{inj}}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)&\text{otherwise}.\end{cases}\end{split}$ (4)
@@ 2.1.2 RPM flow
As a second setup, we consider the rotated potential mixing (RPM) flow Lester et al. (2009). The particle movement is produced using one source and one sink in a two-dimensional plane, operating nonstop and simultaneously. They are placed opposite each other on the edge of the unit circle around zero, which is our domain. The initial locations are $\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ for the source and $\left(\begin{smallmatrix}0\\ -1\end{smallmatrix}\right)$ for the sink. After a time interval of length $\tau$ the source and the sink are rotated around the origin by the angle $\Theta$. Particles that are extracted by the sink are instantaneously injected by the source. Fixing source and sink at their initial positions, we can describe the flow using the Lagrange stream function $\psi$
@@
The analytical solution of the Hamiltonian system is derived in Lester et al. (2009). We solve the flow in the reference system with source and sink at ±$\pm\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$. We give a detailed explanation in Appendix B. Focusing first on the positive side of the $x$-axis (i.e., $x>0$), we represent the particle position in terms of the angle to the origin $\theta\in(-\frac{\pi}{2},\frac{\pi}{2})$ and its streamline $\psi\in(0,\frac{\pi}{2}]$. This is convenient, as $\psi$ is constant in time. Using this coordinate system, the advection time of a particle along a streamline until it reaches $\theta=0$, was established in Lester et al. (2009) as
$\begin{split}t_{\text{adv}}(\theta,\psi)=&\csc^{2}(\psi)\\ \times\Biggl{\{}&\cot(\psi)\arctan\left[\frac{\sin(\theta)\cot(\psi)}{\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}}\right]\\ &+\sin(\theta)\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}\\ &-|\cot(\psi)|(\theta+\cos(\theta)\sin(\theta))\Biggr{\}}.\end{split}$ (6)
The expression uses the cosecant defined as $\csc(x)=\frac{1}{\sin(x)}$ and the cotangent given by $\cot(x)=\frac{1}{\tan(x)}$. The advection time yields a negative value for particles downstream of the centerline $\theta=0$. The residence time of a particle on the streamline $\psi$ is hence given by
@@
$\begin{split}t_{\text{adv}}&(\theta(t),\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}=\\ &\left(t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}-t\right)\mod T_{\text{res}}(\psi_{0}).\end{split}$ (8)
Notice the correction of the sign in front of the $t$ on the right-hand side concerning equation (24) of Lester et al. (2009). A more detailed explanation of the derivation of (7) and (8) is given in Appendix B. Solving this equation yields the flow $\Psi_{1}^{t}$ for all particles on the positive side of the $x$-axis. For particles on the negative side, we use the reflection symmetry of $\psi$ with respect to the $y$-axis, yielding the flow $\Psi_{2}^{t}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)=\left(\begin{smallmatrix}-1&0\\ 0&1\end{smallmatrix}\right)\Psi_{1}^{t}\left(\begin{smallmatrix}-x\\ y\end{smallmatrix}\right)$.
The derivation in Lester et al. (2009) is missing the description for particles on the $y$-axis (i.e., $x=0$) where the usage of $(\theta,\psi)$ coordinates is insufficient. Alternatively, we directly solve (1)
$\begin{matrix}\dfrac{dx}{dt}=\dfrac{4xy}{(y^{2}+x^{2}-1)^{2}+4x^{2}}\\[9.95845pt] \dfrac{dy}{dt}=\dfrac{-2(x^{2}-y^{2}+1)}{(x^{2}+y^{2})^{2}+2x^{2}-2y^{2}+1}\end{matrix}\quad\begin{pmatrix}x(0)\\ y(0)\end{pmatrix}=\begin{pmatrix}0\\ y_{0}\end{pmatrix}$ (9)
@@
$\begin{split}\begin{pmatrix}x(n\tau)\\ y(n\tau)\end{pmatrix}&=R_{n-1}\Psi^{\tau}R_{1-n}...R_{1}\Psi^{\tau}R_{-1}R_{0}\Psi^{\tau}R_{0}\begin{pmatrix}x_{0}\\ y_{0}\end{pmatrix}\\ &=R_{n}\left(R_{-1}\Psi^{\tau}\right)^{n}\begin{pmatrix}x_{0}\\ y_{0}\end{pmatrix}.\end{split}$

### v2

@@ Abstract
Lagrangian particle-tracking methods are particularly suitable to study solute transport in velocity fields displaying chaotic advection. They can accurately resolve stretching and folding processes, the increase in the solute-solvent interface available for diffusion as well as Kolmogorov–Arnold–Moser (KAM) islands, non-mixing regions that limit the chaotic area in the domain and, thereby, the mixing enhancement. However, they also display limitations due to the finite number of discrete particles, particularly if we are interested in the quantification of mixing processes, which require an accurate description of the particle density or concentration gradients. In this work, we use the dilution index to quantify the temporal increase in mixing of a solute within its solvent. We introduce a new approach to select a suitable grid size for the approximation of the density function, motivated by the theory of representative elementary volumes. It preserves the central feature of the dilution index, which is monotonically increasing in time, highlighting the importance of a suitable choice for the grid size in the dilution index approximation. We use this approach to demonstrate the mixing enhancement for two chaotic injection-extraction systems that exhibit chaotic structures: a source-sink dipole and a rotated potential mixing. Using our new approach, we assess the choice of design parameters of the injection-extraction systems to effectively engineer chaotic mixing. We demonstrate the important role of diffusion in filling the KAM islands and reaching complete mixing and, consequently, the importance of avoiding numerical diffusion, which often affects Eulerian methods applied on the advection-diffusion equation.
keywords:
Mixing enhancement , Chaotic advection , Flows in porous media , Dilution index , Representative elementary volume (REV) , Lagrangian particle tracking
@@ 1 Introduction
The natural mixing processes in groundwater and microfluidic systems occurring by molecular diffusion are too slow for many applications [47, 48, 38]. However, advection can boost mixing processes between solute and solvent. This is particularly evident when the flow leads to deformations of the plume, as with shear flow, focusing and defocusing flow, twisting flow [38, 34], or helical flow [57, 58]. It is important to note that while turbulent flows can enhance mixing [55, 18], they are infeasible in certain situations, such as laminar flow in porous media. This limitation underscores the need for alternative methods. One of such alternatives is chaotic advection, a method that has shown potential to enhance fluid mixing [48, 28, 2] and can also occur in flows at low Reynolds numbers [3]. Due to the stretching and folding, the surface between solute and solvent increases, and concentration gradients steepen, boosting the diffusion appearing over this surface [33, 8, 56]. The process of engineering chaotic flow in a system, which requires increased mixing, is called ”design for chaos” [48, 28].
@@
Mixers within laminar flow regimes have a wide range of applications. Beginning with the topic of groundwater remediation, it is shown theoretically [27], numerically [53] and experimentally [43, 16, 22] that the flow through porous media already exhibits chaotic trajectories at the pore scale. At the Darcy scale, the work of [5] gives theoretical evidence that engineered chaotic advection using randomly oscillating wells enhances groundwater remediation if successfully implemented into groundwater flow systems. In practice, an implementation was achieved in [11] where the rotated potential mixing (RPM) flow was engineered at field scale, leading to improved lateral tracer spreading. Moreover, for applications concerning reactive transport the work of [35] highlights that mixing serves as a limiting factor for reactions in diffusion-limited systems, indicating that chaotic mixing can be utilized in a variety of contexts involving reactions. Another application is in microfluidics, where many processes are constrained by incomplete mixing in a laminar flow regime of low Reynolds numbers and high Péclet numbers [1, 54]. There are two main approaches to engineering mixing in these applications: passive and active. A range of passive mixing designs, such as flow-through mixers [48, 28, 42] which are based on a given channel geometry, can be used to increase mixing involving sensitive species [1]. These mixers find application in the synthesis of biomolecules, drug delivery, and diagnostic testing [51]. In active mixers, including microstirrers, acoustic mixers, and flow pulsation [54], chaotic advection is designed using external perturbation sources [1]. A successful implementation of active mixing using pulsating flow was achieved in [48] for enhancing gene expression profiling. Further applications of mixers range from food processing [30] and heat transfer devices [1, 14] to the understanding of chaotic flow through blood vessels [40], showing that chaotic advection and the thereby increased degree of dilution is of high interest.
@@
Modeling transport processes requires solving the classical advection-diffusion equation [20, 6]. Besides the more classical Eulerian methods, like finite differences, finite elements, or finite volumes, one can approximate the solution to this equation using Lagrangian particle tracking. This approach is widely applied in the field of porous media for both conservative [49, 24, 23, 46] and reactive tracers [50, 34, 37, 35, 31, 39]. In contrast to the Eulerian framework, the Lagrangian approach does not suffer from numerical dispersion or artificial oscillations, which requires the Eulerian approach to work with small grid cells and time steps for advection-dominated flows [39, 41, 59]. For conservative transport, the particle tracking approach also has the additional advantage of built-in mass conservation [39]. For chaotic advection setups, the Lagrangian perspective reveals more information about the fractal plume structure [5] and is therefore used for this type of flow, among others, in [4, 5]. The passive advection of a particle is represented by the advection equation, an ordinary differential equation [3], while the diffusive process is added in form of a random walk [5, 15]. For two-dimensional incompressible flows, the advection equation is defined by a Hamiltonian system [7, 3, 2, 10]. The analysis of the resulting advective flow reveals basic properties. Of interest may be Lyapunov exponents, as they show the strength of the stretching and thereby quantify the extent of chaotic flow structures. For volume-preserving flow, we will also find fixed points that are either elliptic or hyperbolic. They alternate on a chain of periodic points and thereby form structures known as island chains [10]. If the flow is chaotic, we may observe the occurrence of Kolmogorov–Arnold–Moser (KAM) islands, which are nonmixing regions arising around the elliptic points in the flow [26, 10]. In the remaining chaotic region, the flow is guided by homoclinic or heteroclinic tangles of the manifolds belonging to unstable stagnation points, leading to exponential fast deformations [10]. Since KAM islands consist of quasi-periodic points [29], particles can only enter them by diffusion. Hence, they limit the chaotic area in the domain and decrease mixing [10].
@@
The goal of stretching and folding is to enlarge the interface between solute and solvent, thereby increasing diffusive fluxes over this interface [5, 32] and enhancing mixing processes. An approach to quantify the effect of mixing is the computation of the dilution index, an entropy-related approach to measure dilution [20]. The term dilution refers to the increase in volume of the fluid occupied by a solute. If the solute concentration follows the advection diffusion equation, this volume increases monotonically in time, hence one of the key properties of the dilution index is that it also increases monotonically in time [20]. For incompressible flow in a non-diffusive system, the volume of a conservative solute and hence also its dilution is constant in time, showing that the diffusion limits the increase in dilution. This measure, applied on results of particle tracking simulations, has a strong dependence on the grid size. It is similar to the problem of estimating the concentration distribution from discrete particle locations that was also identified by [37]. For a non suitable grid the approximation of the dilution index may not show monotonic growth in time, as we see in [19].
@@ 2.1 Chaotic advection setups
In this paper, we consider flow in two dimensions. Hence, to successfully engineer chaos, the advection equation must be unsteady [33, 48, 3]. By following a large number of tracer particles inside the system, we approximate the spatial concentration distribution of the plume over time.
@@ 2.1.1 Pulsed source-sink system
The first setup under consideration is called the pulsed source-sink system introduced in [17] and later also addressed in [44, 48, 45]. We consider a two-dimensional plane with one sink and one source at arbitrary locations (see Fig. 1). During the experiment, the sink and the source are operated alternatively for a time interval of length $\tau$, meaning only one is switched on at a time. The particles extracted by the sink are reinjected by the source in the next stroke using the first-out-first-in procedure [48]. This can be realized physically by collecting the extracted particle in a tube beneath the flow plane during the stoke of the sink. Prior to the stroke of the source the tube is flipped over and connected to the source where the fluid is then reinjected into the plane [17].
Figure 1: Two-dimensional plane with one sink and one source at $\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. All particles with a distance to the sink smaller than $\Lambda$ are extracted during the next stroke of the sink. Within the following stroke of the source, all of them are reinjected and occupy the circle around the source of radius $\Lambda$.
In general, the flow of an incompressible fluid in a non-deformable medium is described by a Lagrange stream function [7, 3, 2, 10]. The existence of a Lagrange stream function $\psi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ implies that the associated flow can be described using the Hamiltonian system [44, 12]
@@
For a point source or sink at location $\left(\begin{smallmatrix}\bar{x}\\ \bar{y}\end{smallmatrix}\right)$, $\psi$ is given in [45] by
@@
The analytical solution of the flow defined in (1) is derived in [45, 17]. We give a detailed derivation in Appendix A. Arguing that each positioning of sink and source can be transformed into an equivalent configuration with sink and source at $\pm\left(\begin{smallmatrix}a\\ 0\end{smallmatrix}\right)$ and by multiplying the length scale by $1/a$, we assume sink and source at locations $\pm\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$. The flow is parametrized by a combination of $\tau$, $Q$, and $a$ in the definition of
@@
It holds that all particles within the circle of radius $\Lambda$ around the sink are extracted during the operation of the sink for time $\tau$. This circle is depicted as $B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$ and also shown in Fig. 1. For the extraction-reinjection procedure, the authors of [17] derive the map $M_{\text{inj}}:B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\rightarrow\mathbb{R}^{2}$, taking into consideration the first-out-first-in reinjection protocol and an injection angle of $\pi-\theta$ given the extraction angle $\theta$. The map is given by
@@
$M\begin{pmatrix}x\\ y\end{pmatrix}=\begin{cases}\left(\Psi_{+}^{\tau}\circ\Psi_{-}^{\tau}\right)\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)&\text{for }\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)\in\mathbb{R}^{2}/B_{\Lambda}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)\\ M_{\text{inj}}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)&\text{otherwise}.\end{cases}$ (4)
@@ 2.1.2 RPM flow
As a second setup, we consider the rotated potential mixing (RPM) flow [26]. The particle movement is produced using one source and one sink in a two-dimensional plane, operating nonstop and simultaneously. They are placed opposite each other on the edge of the unit circle around zero, which is our domain. The initial locations are $\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$ for the source and $\left(\begin{smallmatrix}0\\ -1\end{smallmatrix}\right)$ for the sink. After a time interval of length $\tau$ the source and the sink are rotated around the origin by the angle $\Theta$. Particles that are extracted by the sink are instantaneously injected by the source. Fixing source and sink at their initial positions, we can describe the flow using the Lagrange stream function $\psi$
@@
The analytical solution of the Hamiltonian system is derived in [26]. We solve the flow in the reference system with source and sink at $\pm\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$. We give a detailed explanation in Appendix B. Focusing first on the positive side of the $x$-axis (i.e., $x>0$), we represent the particle position in terms of the angle to the origin $\theta\in(-\pi/2,\pi/2)$ and its streamline $\psi\in(0,\pi/2]$. This is convenient, as $\psi$ is constant in time. Using this coordinate system, the advection time of a particle along a streamline until it reaches $\theta=0$, was established in [26] as
$\begin{split}t_{\text{adv}}(\theta,\psi)=\csc^{2}(\psi)\times\Biggl\{&\cot(\psi)\arctan\left[\frac{\sin(\theta)\cot(\psi)}{\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}}\right]\\ &+\sin(\theta)\sqrt{1+\cos^{2}(\theta)\cot^{2}(\psi)}-|\cot(\psi)|(\theta+\cos(\theta)\sin(\theta))\Biggr\}.\end{split}$ (6)
The advection time yields a negative value for particles downstream of the centerline $\theta=0$. The residence time of a particle on the streamline $\psi$ is hence given by
@@
$\begin{split}t_{\text{adv}}(\theta(t),\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}=\left(t_{\text{adv}}(\theta_{0},\psi_{0})+\frac{T_{\text{res}}(\psi_{0})}{2}-t\right)\mod T_{\text{res}}(\psi_{0}).\end{split}$ (8)
Notice the correction of the sign in front of the $t$ on the right-hand side concerning equation (24) of [26]. A more detailed explanation of the derivation of (7) and (8) is given in Appendix B. Solving this equation yields the flow $\Psi_{1}^{t}$ for all particles on the positive side of the $x$-axis. For particles on the negative side, we use the reflection symmetry of $\psi$ with respect to the $y$-axis, yielding the flow $\Psi_{2}^{t}\left(\begin{smallmatrix}x\\ y\end{smallmatrix}\right)=\left(\begin{smallmatrix}-1&0\\ 0&1\end{smallmatrix}\right)\Psi_{1}^{t}\left(\begin{smallmatrix}-x\\ y\end{smallmatrix}\right)$.
The derivation in [26] is missing the description for particles on the $y$-axis (i.e., $x=0$) where the usage of $(\theta,\psi)$ coordinates is insufficient. Alternatively, we directly solve (1)
$\dfrac{dx}{dt}=\dfrac{4xy}{(y^{2}+x^{2}-1)^{2}+4x^{2}},\qquad\dfrac{dy}{dt}=\dfrac{-2(x^{2}-y^{2}+1)}{(x^{2}+y^{2})^{2}+2x^{2}-2y^{2}+1},\qquad\begin{pmatrix}x(0)\\ y(0)\end{pmatrix}=\begin{pmatrix}0\\ y_{0}\end{pmatrix}$ (9)
@@
$\begin{pmatrix}x(n\tau)\\ y(n\tau)\end{pmatrix}=R_{n-1}\Psi^{\tau}R_{1-n}...R_{1}\Psi^{\tau}R_{-1}R_{0}\Psi^{\tau}R_{0}\begin{pmatrix}x_{0}\\ y_{0}\end{pmatrix}=R_{n}\left(R_{-1}\Psi^{\tau}\right)^{n}\begin{pmatrix}x_{0}\\ y_{0}\end{pmatrix}$.